OpenStudy (anonymous):

r = 8sin(Theta). Write an equation in rectangular coordinates.

OpenStudy (irishboy123):

what are x and y in polar?!?!

OpenStudy (anonymous):

"Transform the polar equation to an equation in rectangular coordinates" is all it says ...

OpenStudy (anonymous):

but everyone is telling me to use this formula: x = rcosθ y = rsinθ x² + y² = r²

OpenStudy (irishboy123):

yes that's right so \(r = 8 \sin( \theta)\) is the same as \(\sqrt{x^2 + y^2} = 8\dfrac{ y}{\sqrt{x^2 + y^2}}\)

OpenStudy (anonymous):

Yes but I am not sure how to solve this..

OpenStudy (irishboy123):

multiply each side by \(\sqrt{x^2 + y^2}\)

OpenStudy (irishboy123):

\[\sqrt{x^2 + y^2} * \sqrt{x^2 + y^2} = 8\dfrac{ y}{\sqrt{x^2 + y^2}}*\sqrt{x^2 + y^2} \]

OpenStudy (anonymous):

??

OpenStudy (irishboy123):

\[\sqrt{x^2 + y^2} * \sqrt{x^2 + y^2} = x^2 + y^2 \] right?!

OpenStudy (irishboy123):

and \[\dfrac{ y}{\sqrt{x^2 + y^2}}*\sqrt{x^2 + y^2} = y\] yes?